OpenStudy (taepoohwae):

the cost of driving a car includes both fixed costs and mileage costs. assume that it costs $162.10 per month for insurance and car payments and $0.26 per mile gasoline, oil, and routine maintenance. Find the values for m and b so that y=mx+b models the monthly cost of driving the car x miles. what dos the value of b represent?

OpenStudy (anonymous):

y = mx + b y is total cost m is 0.26 (cost per mile, gas,oil, and maintenance) x is number of miles b = 162.10 (cost per month for insurance and car payments y = 0.26x + 162.10 <===here is the equation Do you understand ?

OpenStudy (taepoohwae):

I'm new to this. I really don'tr know where to begin in order to solve.

OpenStudy (anonymous):

Look at what I did and tell me what you don't understand ?

OpenStudy (taepoohwae):

Not sure of the steps to follow to find the answer to m or b.

OpenStudy (taepoohwae):

Are you saying that the answer for m is 0.26 and b is 162.10

OpenStudy (anonymous):

The question asks for the value of mx and b. m = 0.26 with x being undefined number of miles. b = monthly cost of the car.

OpenStudy (anonymous):

let me rephrase that....y = the total cost (gas, car payment,maintenance, etc. b is just the cost of the car payment and insurance. m is the cost per mile, gas, oil and stuff like that (.26) x being the number of miles (which is not given)

OpenStudy (taepoohwae):

Thanks for trying... I still don't know how to solve for m or b.

OpenStudy (anonymous):

@Hero ...can you help me explain this ?

OpenStudy (anonymous):

hold on taepoohwae.....Hero is better at explaining then I am. He will be here soon

OpenStudy (anonymous):

@Hero .....help

OpenStudy (anonymous):

@ParthKohli ....can you help me explain this ?

OpenStudy (anonymous):

taepoohwae doesn't seem to understand ...I am not really good at explaining

Parth (parthkohli):

Start with ignoring the \(y = mx + b\) thingy. It's of no use. So you need \(0.26\) for one mile. Then you need \(0.26 \times \text{number of miles driven}\) throughout the month. Assume that \(m\) is the number of miles you drive, then \(0.26m\) is the money you need. You write \(y = 0.26m\) where \(y\) is the cost. STOP RIGHT THERE! You need \(162.10\) for maintenance, so it'd always be there added to the costs like a ninja. That'd make \(0.26m + 162.10\) as the total money you need

OpenStudy (anonymous):

Oh...your good, you should be a teacher

Parth (parthkohli):

Actually let's just say that the number of miles is \(x\). Then you have \(y = \color{#C00}{m}x + \color{#AE9}{b}\)

Parth (parthkohli):

\[y = \color{#C00}{m}x + \color{#AE5}{b}\]\[y = \color{#C00}{0.26}x + \color{#AE3}{162.10}\]

OpenStudy (anonymous):

maybe taepoohwae will understand now....thx ParthKohli

OpenStudy (taepoohwae):

thanks!

OpenStudy (anonymous):

your welcome :)